For n = 121, sample mean = 96 and a known population standard deviation = 14, construct a 95% confidence interval for the population mean.
Answer #1

Solution :

Given that,

Sample size = n = 121

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.96 * (14 / \sqrt 121)

Margin of error = E = 2.5

At 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

96 - 2.5 < \mu < 96 + 2.5

93.5 < \mu < 98.5

(93.5, 98.5)